# Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage

### v3

@@ 5 Experimental Result
Also, as shown in Table VII, LLM and some indicators of downstream tasks, such as BLUE [40] and ROUGH [41], used to evaluate translation tasks, also uses abbreviations in some tables for easier reading. For the experimental metrics, the best metric is bolded and the suboptimal metric is underlined.
| Category | Full Name | Abbreviation |
| LLM | GPT-3.5-turbo | G-35T |
| GPT-4o | G-4o | |
| GPT-o1-mini | Go1-m | |
| DeepSeek-R1 | DK-r1 | |
| LlaMA-3.18B | L-318B | |
| Qwen-2.5 | Q-2.5 | |
| Sun-Shine | S-S | |
| Evaluation Metrics | BLEU-4 | B-4 |
| ROUGE-1 | R-1 | |
| ROUGE-2 | R-2 | |
| ROUGE-l | R-l |
5.1 Insights from Performance Results
Ti-MMLU: as shown in Fig. 9, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
For Response Rate, as shown in Table VIII, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1. Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training benchmark.
@@ 5.1 Insights from Performance Results
For Conditional Accuracy, as shown in Table IX, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead. Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
@@
For Accuracy, as shown in Table X, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
@@
This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient. Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
As shown in Fig. 10, we also compiled statistics on various indicators of Sun-Shine in each sub-branch. Whether it is the best or the second best indicator, Sun-Shine maintains a large advantage in total, whether compared with DeepSeek-R1 or Chat-4o. Especially in the sub-category, “Other”, Sun-Shine obtained the best indicators of all accuracy. The remaining sub-categories also outperform DeepSeek-r1 overall. Only in the indicator ”Conditional Accuracy”, Sun-Shine is slightly lower than the latest DeepSeek-R1.
Also, as shown in Table XI, we also calculated the average of different metrics for different large models on different categories. Sun-Shine is either first or second in all categories and all indicators. Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
| LLM | Response Rate | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.5507 | 0.5588 | 0.5377 | 0.5481 | 0.5498 |
| GPT-3.5-turbo | 0.0687 | 0.0741 | 0.0748 | 0.0806 | 0.0801 |
| GPT-4o | 0.4932 | 0.5133 | 0.5107 | 0.4915 | 0.4957 |
| GPT-o1-mini | 0.2050 | 0.2340 | 0.2333 | 0.2231 | 0.2241 |
| Llama-3.18B | 0.1907 | 0.1998 | 0.2056 | 0.2082 | 0.2104 |
| Qwen-2.5 | 0.2563 | 0.2703 | 0.2736 | 0.2441 | 0.2419 |
| DeepSeek-R1 | 0.4206 | 0.4069 | 0.3884 | 0.3580 | 0.3482 |
| LLM | Conditional Accuracy | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.3208 | 0.3367 | 0.3477 | 0.3457 | 0.3489 |
| GPT-3.5-turbo | 0.2489 | 0.2605 | 0.2656 | 0.2710 | 0.2731 |
| GPT-4o | 0.2845 | 0.2871 | 0.2846 | 0.2746 | 0.2775 |
| GPT-o1-mini | 0.2636 | 0.2700 | 0.2716 | 0.2742 | 0.2728 |
| Llama-3.18B | 0.2510 | 0.2704 | 0.2662 | 0.2620 | 0.2597 |
| Qwen-2.5 | 0.2656 | 0.2655 | 0.2710 | 0.2866 | 0.2900 |
| DeepSeek-R1 | 0.3209 | 0.3482 | 0.3524 | 0.3584 | 0.3573 |
| LLM | Accuracy | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.1888 | 0.1742 | 0.1683 | 0.1614 | 0.1618 |
| GPT-3.5-turbo | 0.0182 | 0.0201 | 0.0205 | 0.0277 | 0.0265 |
| GPT-4o | 0.1594 | 0.1676 | 0.1634 | 0.1515 | 0.1535 |
| GPT-o1-mini | 0.0475 | 0.0584 | 0.0598 | 0.0603 | 0.0606 |
| Llama-3.18B | 0.0656 | 0.0628 | 0.0723 | 0.0662 | 0.0663 |
| Qwen-2.5 | 0.0691 | 0.0723 | 0.0744 | 0.0702 | 0.0709 |
| DeepSeek-R1 | 0.1863 | 0.1819 | 0.1705 | 0.1562 | 0.1534 |
Ti-Safetybench: as shown in Fig. 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
As shown in Table XII, it presents a comprehensive comparison of leading LLMs on the Ti-Safetybench, evaluating their performance across key safety-related categories. Notably, Sun-Shine outperforms all baseline models under both DA and CCA tasks. In the DA setting, Sun-Shine achieves the highest exact match (EM) score (0.292) and leads in IA, OFF, and UB, while maintaining a significantly higher filter rate (59.20%) than GPT-4o (38.59%) and others. In the CCA setting, Sun-Shine demonstrates even stronger results, ranking first across nearly all categories, including EM (0.483), IA (0.455), mental health (MH, 0.507), and UB (0.466). These results highlight Sun-Shine’s superior ability to generate safe, accurate, and instruction-aligned responses, setting a new benchmark in safety-aware language modeling.
| LLM | Category | ||||||||
| Filter | Invaild | EM | IA | MH | OFF | PH | PP | UB | |
| Sun-Shine | DA | 59.20% | 0.292 | 0.379 | 0.204 | 0.255 | 0.258 | 0.307 | 0.372 |
| GPT-4o | 38.59% | 0.144 | 0.357 | 0.241 | 0.199 | 0.246 | 0.296 | 0.358 | |
| Llama-3.18B | 74.75% | 0.105 | 0.088 | 0.101 | 0.097 | 0.094 | 0.099 | 0.109 | |
| GPT-3.5-turbo | 73.42% | 0.085 | 0.073 | 0.087 | 0.086 | 0.097 | 0.106 | 0.123 | |
| Qwen-2.5 | 49.50% | 0.199 | 0.118 | 0.164 | 0.198 | 0.150 | 0.159 | 0.371 | |
| GPT-o1-mini | 71.20% | 0.109 | 0.070 | 0.095 | 0.111 | 0.097 | 0.095 | 0.170 | |
| DeepSeek-R1 | 51.14% | 0.287 | 0.373 | 0.232 | 0.250 | 0.297 | 0.309 | 0.371 | |
| Sun-Shine | CCA | 16.60% | 0.483 | 0.455 | 0.507 | 0.448 | 0.509 | 0.336 | 0.466 |
| GPT-4o | 34.55% | 0.478 | 0.429 | 0.429 | 0.287 | 0.255 | 0.387 | 0.401 | |
| Llama-3.18B | 69.72% | 0.130 | 0.101 | 0.123 | 0.124 | 0.119 | 0.115 | 0.126 | |
| GPT-3.5-turbo | 67.75% | 0.108 | 0.082 | 0.103 | 0.125 | 0.114 | 0.118 | 0.160 | |
| Qwen-2.5 | 16.61% | 0.448 | 0.413 | 0.427 | 0.289 | 0.259 | 0.385 | 0.403 | |
| GPT-o1-mini | 61.69% | 0.254 | 0.099 | 0.134 | 0.163 | 0.135 | 0.117 | 0.231 | |
| DeepSeek-R1 | 28.14% | 0.558 | 0.457 | 0.500 | 0.429 | 0.511 | 0.339 | 0.457 | |
Overall: Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
5.2 Sun-Shine for Tibetan Culture
As shown in Fig. 12, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
As shown in Fig. 13(a), Fig. 13(b), and Fig. 13(c), we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
Sun-Shine generates outputs that are notably neutral, emotionally grounded, and culturally restrained across all three languages. The Tibetan response is respectful and sincere, avoiding ornate metaphors or religious references. In Chinese, the model articulates gratitude with clarity and warmth, emphasizing familial support without exaggeration. The English version maintains a reflective tone, expressing appreciation without invoking religious or hyperbolic imagery. This demonstrates Sun-Shine’s alignment with real-world communication norms and its ability to generalize across linguistic and cultural contexts.
In contrast, GPT-4o exhibits a distinctly elevated and ceremonial tone. Its Tibetan and Chinese outputs are imbued with grandiloquent phrases and spiritual imagery. For example, the model refers to the father as a “beacon of happiness” and ”born on the edge of happiness,” which, while poetic, may not align with typical expressions of filial appreciation. The English version similarly includes metaphor-heavy language, such as “the victory banner is my prayer,” which resembles religious scripture more than natural expression. This suggests a stylistic bias toward exalted, sometimes deific, portrayals.
DeepSeek-R1, on the other hand, appears to have an even stronger inclination toward religious and cultural iconography. It references “Mañjuśrī’s Father” and constructs its Tibetan output with vocabulary and structure closely mirroring classical Buddhist literature. The Chinese response is elegant yet heavily stylized, and the English output reads like a translated verse from scripture rather than a naturally composed poem. While aesthetically refined, this stylistic preference may reflect overfitting to religious texts in training data, limiting the model’s adaptability to general-purpose use.
In summary, Sun-Shine’s responses are more balanced and less prone to embellishment, demonstrating cultural sensitivity and instruction alignment. GPT-4o and DeepSeek-R1, while fluent and creative, tend to overgenerate in stylistically loaded contexts, risking a mismatch with the intended pragmatic tone. This comparison underscores Sun-Shine’s strength in producing appropriate, grounded, and respectful language across multiple languages and cultural domains.
5.3 Downstream Task
Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
| LLM | A | B | C | D |
| Sun-Shine | 63 | 61 | 44 | 42 |
| GPT-4o | 45 | 47 | 53 | 51 |
| DeepSeek-R1 | 59 | 56 | 67 | 66 |
| Qwen-2.5 | 41 | 43 | 53 | 54 |
As shown in Table XIII, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our benchmark, TIB-STC, which includes a richer collection of ancient Tibetan texts.
Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU [40] and ROUGE [41] are used to evaluate this task. As shown in Table XIV, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 12 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
| LLM | B-4 | R-1 | R-2 | R-l |
| Sun-Shine | 42.16 | 65.78 | 48.91 | 51.58 |
| GPT-4o | 38.85 | 61.12 | 44.33 | 47.02 |
| DeepSeek-R1 | 41.85 | 65.49 | 47.73 | 51.04 |
| Qwen-2.5 | 35.20 | 58.59 | 41.25 | 44.73 |
6 Conclusion
In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC benchmark, marking significant progress in Tibetan language processing. Sun-Shine, trained on the comprehensive TIB-STC benchmark, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios. TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data. Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan. Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our benchmark, TIB-STC, in providing a high-quality ancient Tibetan corpus. This work advances low-resource language models and underscores the importance of culturally specific benchmarks.
Acknowledgment
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.

### v4

@@ Introduction
Detailed experimental settings and evaluation metrics for each model are provided in Table 4. Specifically, Qwen-2.5 and DeepSeek-R1 were tested via their official APIs, while LLaMA-3 18B was accessed through the LLaMA-API platform. All experimental configurations will be fully documented in the final version to ensure transparency and reproducibility.
Insights from Performance Results
Ti-MMLU
As shown in Figure 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs. Among them, Sun-Shine is slightly ahead of these. For the Response Rate, Sun-Shine and GPT-4o achieve the highest Response Rate, with Sun-Shine leading across all sub-branches,shown in Figure 8, demonstrating superior responsiveness to Tibetan prompts enabled by its domain-specific training. For Conditional Accuracy, Sun-Shine and DeepSeek-R1 perform best, with the latter holding a slight advantage. Both models maintain consistently high scores, reflecting strong general knowledge capabilities. In terms of overall Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas. This underscores that Sun-Shine’s strong Tibetan understanding stems not only from model design but also from effective use of the TIB-STC dataset and targeted training, allowing an 8B model to surpass the 671B MoE DeepSeek-R1 in key tasks.
Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. For more details, please refer to supplementary materials.
Ti-Safetybench
as shown in Figure 9, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
Sun-Shine for Tibetan Culture
As shown in Figure 10, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
As shown in Figure 11, we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English. Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics. In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o. Overall, Sun-Shine surpasses DeepSeek-R1 and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
Downstream Task
Translation
we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU (Papineni et al. 2002) and ROUGE (Lin 2004) are used to evaluate this task. As shown in Table 5, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 10 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
| LLM | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-l |
| Sun-Shine | 42.16 | 65.78 | 48.91 | 51.58 |
| GPT-4o | 38.85 | 61.12 | 44.33 | 47.02 |
| DeepSeek-R1 | 41.85 | 65.49 | 47.73 | 51.04 |
| Qwen-2.5 | 35.20 | 58.59 | 41.25 | 44.73 |
Tibetan Language Proficiency Test
we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system). As shown in Table 6, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
| LLM | A | B | C | D |
| Sun-Shine | 63 | 61 | 44 | 42 |
| GPT-4o | 45 | 47 | 53 | 51 |
| DeepSeek-R1 | 59 | 56 | 67 | 66 |
| Qwen-2.5 | 41 | 43 | 53 | 54 |
Conclusion
In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC benchmark and the Sun-Shine large language model. Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research. Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models. Leveraging this benchmark, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression. Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment. These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded benchmarks in advancing low-resource language technologies.
Future Work
As part of our ongoing efforts, we have launched Sun-Shine 2.0, an upgraded version of our Tibetan foundation model with enhanced reasoning capabilities. In parallel, we are curating an extended Tibetan corpus to further improve linguistic and cultural coverage.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62276055 and 62406062, in part by the Sichuan Science and Technology Program under Grant 2023YFG0288, in part by the Natural Science Foundation of Sichuan Province under Grant 2024NSFSC1476, in part by the National Science and Technology Major Project under Grant 2022ZD0116100, and in part by the Sichuan Provincial Major Science and Technology Project under Grant 2024ZDZX0012.
Appendix A Supplementary Material
Note: S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1.
Details of Response Rate
@@ Appendix A Supplementary Material
Details of Conditional Accuracy
@@
Details of Accuracy